# How important is the dielectric constant in water modeling? Evaluation of the performance of the TIP4P/ force field and its compatibility with the Joung-Cheetham NaCl model

### v1

@@ 5.1 Joung-Cheetham NaCl model
5.2 Density as a function of molality
@@
5.3 Temperature of maximum density
@@
The TMD is quite sensitive to the addition of salts. Results for the density of JC salts of different concentrations dissolved in TIP4P/ water are shown in Figure 7-b. From these plots, the temperature of maximum density (TMD) has been estimated as shown in Table 5, which also compares results for JC ions in TIP4P/2005 water obtained previously 86, 88. It is clear that both models reproduce experimental results reasonably well at all considered NaCl concentrations, despite the use of elementary charges for the ions.
| m (mol/kg) | Exp. | JC-TIP4P/2005 | JC-TIP4P/ |
| 0 | 277.1 | 278 | 276 |
| 0.5 | 270.5 | 267 | 269 |
| 1.0 | 263.5 | 259 | 258/261 |
5.4 Dielectric constant as a function of molality
@@
Dielectric properties are also quite sensitive to salt concentration, so it is interesting to test how this dependence is described by TIP4P/, which already exhibits a good dielectric constant for pure water. Figure 8-a. shows that TIP4P/ can predict very accurately a large drop of the dielectric constant with salt concentration, providing quantitative results for up to two molal concentrations. Results for larger concentrations deteriorate, but remain quite reasonable up to the measured concentrations of about six molal.
@@
A similar performance for the TIP4P/2005 model can obviously not be expected, since the dielectric constant is predicted already 25% too low for pure water. However, it is interesting to note that the strong decrease of the dielectric constant is properly predicted by both models. To see this, Figure 8-b displays the relative drop of the dielectric constant, . In this scale, both models behave similarly. TIP4P/ predicts to sharp a drop of the relative dielectric constant, while TIP4P/2005 appears to predict values somewhat larger than found in experiment.
5.5 Viscosity as a function of molality
@@
As a final check, we now test the role of salt concentration on the viscosity of the solutions in Figure 9. Unfortunately, in this case the results for the JC ions in combination with the TIP4P/ are very disappointing, with predicted viscosities that are far too large in the entire concentration range studied. This has been observed previously for dynamic properties of electrolyte solutions of rigid point charge solvents. The large charge of the ions leads to trapping of the water molecules, which have great difficulty escaping from the solvation cage for lack of local charge fluctuations in the rigid point charge models 25. This explains why the Madrid-2019 force field, with ionic charges scaled to 0.85 times the elementary charge performs significantly better. Blazquez et al. 25 have shown that the discrepancy can be removed by further scaling the charges to even smaller values (equal to ), but unfortunately, this is not satisfactory either, as other properties are then not accurately predicted.
5.6 Melt and lattice properties
@@
Since the JC model was parametrized with the NaCl crystal and melt as a target, the properties of the melt and the solid at the melting point are also rather accurately predicted, as shown in Table 6. On the other hand, the Madrid-2019 force field is somewhat less accurate, as this set of properties were sacrificed for better performance of solution properties.
Overall, this means that TIP4P/ appears as a rather robust model for the study of solution properties, as it provides a rather fair account of solution properties for models parametrized for the pure salt properties. This makes the model transferable to a significant extent, and is expected to provide an accurate account of the phase coexistence of aqueous electrolytes.
@@
6 Conclusions
In this work we have considered the role of the dielectric constant as a parametrization target for accurate empirical force fields. Based on previous results, including some of our own work, we had expected that an accurate account of the dielectric constant of simple point charge models could be at the cost of compromising the accuracy of a large number of properties 69, 27, 41. A dramatic example is the case of the SPC/E model, which predicts a much better dielectric constant than most TIP4P models, but is unable to provide a qualitatively correct description of the solid phases 13. A more recent example is the work of Sedano et al., who showed that OPC, a well known model parametrized with the dielectric constant, performs significantly less well than TIP4P/2005 as based on the Vega-Abascal benchmark 41.
These observations can be rationalized to some extent by the scaled-charge paradigm, which dictates that the nominal point charges of empirical force fields are screened by the electronic polarization of the surrounding medium 17, 18. An implication is that the charges that dictate the force field are scaled versions of the charges that dictate the dipole moment 69, 27. Whence, one expects that targeting the point charges to the dielectric constant will result in too large charges and dipole moments for the force field.
Contrary to these expectations, our results show that the TIP4P/ model, which has been parametrized to predict accurate densities and dielectric constants, can also provide a rather accurate account of many other properties with a similar overall accuracy as the acclaimed TIP4P/2005 model. Indeed, we find that TIP4P/ can predict bulk phase properties (transport properties, thermal expansion coefficients, densities) of both liquid water and ice polymorphs, as well as surface tensions, with an accuracy very similar to TIP4P/2005, while performing very well for dielectric constants over a wide range of pressures and temperatures. On the other hand, TIP4P/2005 provides a better description of phase boundaries, including liquid-vapor and freezing transitions. Overall, these two models appear to perform almost identically on the conventional benchmark 16 while significantly outperforming the OPC force field.
@@ 6 Conclusions
Unfortunately, the good description of solution thermodynamics of TIP4P/ breaks down when one seeks an accurate account of transport properties, a problem that is also shared by the TIP4P/2005 model 25. This points to the limitations of point charge models and the need of addressing the problem of polarization in a physically meaningful way for next generation force fields with enhanced transferability. Current advances in computer architecture might well remedy the computational overhead and make such models competitive in the very near future. In the meantime, TIP4P/2005 and TIP4P/ force fields appear as the two equally valid point charge models for use in routine calculations.
An intriguing question that remains unsolved is which of these two models is a better caricature of the actual charge distribution and local electric field created by a water molecule in the condensed phase. However, because this is actually a collective property that depends on many body interactions, it might simply be an ill-posed question. Be as it may, calculations of the dielectric constant of these models indicate that they predict a rather different Kirkwood factor, and accordingly, significantly different orientation correlations in the liquid phase. Measuring such correlations experimentally appears difficult, but would allow to discriminate between these two seemingly different empirical force fields. As a corollary, solving this problem would also help to assess the validity of charge scaling. However, as long as this is not clarified, the fact that a model mainly targeted for the dielectric constant can perform as well as TIP4P/2005 with improved transferability in solution, casts some doubts on the otherwise appealing charge scaling paradigm.
We would like to thank Eva G. Noya and Enrique Lomba for helpful discussions. ŁB would like to thank the Polish National Agency for Academic Exchange, under Grant No. BPN/BEK/2023/1/00006 Bekker 2023. CDAG acknowledges the predoctoral fellowship ”Contratos Predoctorales de Personal Investigador en Formación” from the Universidad Complutense de Madrid (CT15/23) and additional funding from Ministerio de Ciencia, Innovación y Universidades (MICIU), under grant FPU22/00869. LGM would like to thank financial support from the Agencia Estatal de Investigación (Ministerio de Ciencia y Economía) under grant PID2023-151751NB-I00, funded by MCIN/AEI/10.13039/501100011033. We also benefited from generous allocation of computer time at the Academic Supercomputer Center (CI TASK) in Gdansk.

### v2

@@ V.1 Joung-Cheatham NaCl model
V.2 Density as a function of molality
@@
V.3 Temperature of maximum density
@@
The TMD is quite sensitive to the addition of salts. Results for the density of JC salts of different concentrations dissolved in TIP4P/ water are shown in Figure 7-b. From these plots, the temperature of maximum density (TMD) has been estimated as shown in Table 5, which also compares results for JC ions in TIP4P/2005 water obtained previously Benavides, Aragones, and Vega (2016); Benavides et al. (2017). It is clear that both models reproduce experimental results reasonably well at all considered NaCl concentrations, despite the use of elementary charges for the ions. The same conclusion can be drawn regarding the Madrid-2019 force field. However, the main advantage is that it not only captures the TMD but also reproduces the experimental densities within a wide temperature range which is expected based on the Figure 7-a.
| m (mol/kg) | Exp. | JC-TIP4P/2005 | JC-TIP4P/ | Madrid-2019 |
| 0 | 277.1 | 278 | 276 | 278 |
| 0.5 | 270.5 | 267 | 269 | |
| 1.0 | 263.5 | 259 | 258/261 | 260.7 |
V.4 Solvent dielectric constant as a function of molality
Dielectric properties are also quite sensitive to salt concentration, so it is interesting to test how this dependence is described by TIP4P/, which already exhibits a good dielectric constant for pure water. However, one needs to bear in mind that due to the conducting nature of electrolyte solutions, the dielectric constant measured by Eq. 7, which holds for an insulating media, is not directly accessible from the experiments Hubbard, Colonomos, and Wolynes (1979); Sanchez et al. (2018); Sanchez, Dominguez, and Pizio (2020). In the latter case, the dielectric constant is measured by extrapolation to zero frequency after the removal of the divergent conductive contributions. However, the resulting property peaks up additional dynamical contributions related to cross correlations between the dipole moment and the charge current Caillol, Levesque, and Weis (1986) that are not contained in Eq. 7. Although this contribution appears to be relatively small,Caillol, Levesque, and Weis (1986) a direct comparison with experimental data shown should be taken with some caution.
@@ V.4 Solvent dielectric constant as a function of molality
Figure 8-a. shows that TIP4P/ can predict very accurately a sharp drop of the dielectric constant with salt concentration, providing quantitative results for up to two molal concentrations. Results for larger concentrations deteriorate but remain quite reasonable up to the measured concentrations of about six molal.
@@
A similar performance for the TIP4P/2005 model can obviously not be expected, since the dielectric constant is predicted already 25% too low for pure water. However, it is interesting to note that the strong decrease in the dielectric constant is properly predicted by both models. To see this, Figure 8-b displays the relative drop of the dielectric constant, . In this scale, both models behave similarly. TIP4P/ predicts to sharp a drop of the relative dielectric constant, while TIP4P/2005 appears to predict values somewhat larger than found in the experiment.
V.5 Viscosity as a function of molality
@@
As a final check, we now test the role of salt concentration on the viscosity of the solutions in Figure 9. Unfortunately, in this case, the results for the JC ions in combination with the TIP4P/ are very disappointing, with predicted viscosities that are far too large in the entire concentration range studied. This has been observed previously for the dynamic properties of electrolyte solutions of rigid point charge solvents. The large charge of the ions leads to trapping of the water molecules, which have great difficulty escaping from the solvation cage for lack of local charge fluctuations in the rigid point charge models Blazquez, Conde, and Vega (2023). This explains why the Madrid-2019 force field, with ionic charges scaled to 0.85 times the elementary charge performs significantly better. Blazquez et al. Blazquez, Conde, and Vega (2023) have shown that the discrepancy can be removed by further scaling the charges to even smaller values (equal to ), but unfortunately, this is not satisfactory either, as other properties are then not accurately predicted.
V.6 Melt and lattice properties
@@
Since the JC model was parametrized with the NaCl crystal and melt as a target, the properties of the melt and the solid at the melting point are also rather accurately predicted, as shown in Table 6. On the other hand, the Madrid-2019 force field is somewhat less accurate, as this set of properties was sacrificed for better performance of solution properties.
Overall, this means that TIP4P/ appears as a rather robust model for the study of solution properties, as it provides a rather fair account of solution properties for models parametrized for the pure salt properties. This makes the model transferable to a significant extent and is expected to provide an accurate account of the phase coexistence of aqueous electrolytes.
@@
VI Conclusions
In this work, we have considered the role of the dielectric constant as a parametrization target for accurate empirical force fields. Based on previous results, including some of our own work, we had expected that an accurate account of the dielectric constant of simple point charge models could be at the cost of compromising the accuracy of a large number of properties MacDowell and Vega (2010); Aragones, MacDowell, and Vega (2011a); Sedano, Blazquez, and Vega (2024). A dramatic example is the case of the SPC/E model, which predicts a much better dielectric constant than most TIP4P models, but is unable to provide a qualitatively correct description of the solid phases Sanz et al. (2004). A more recent example is the work of Sedano et al., who showed that OPC, a well-known model parametrized with the dielectric constant, performs significantly less well than TIP4P/2005 as based on the Vega-Abascal benchmark Sedano, Blazquez, and Vega (2024).
These observations can be rationalized to some extent by the scaled-charge paradigm, which dictates that the nominal point charges of empirical force fields are screened by the electronic polarization of the surrounding medium Leontyev and Stuchebrukhov (2009, 2011). An implication is that the charges that dictate the force field are scaled versions of the charges that dictate the dipole moment MacDowell and Vega (2010); Aragones, MacDowell, and Vega (2011a). Whence, one expects that targeting the point charges to the dielectric constant will result in too large charges and dipole moments for the force field.
Contrary to these expectations, our results show that the TIP4P/ model, which has been parametrized to predict accurate densities and dielectric constants, can also provide a rather accurate account of many other properties with similar overall accuracy as the acclaimed TIP4P/2005 model. Indeed, we find that TIP4P/ can predict bulk phase properties (transport properties, thermal expansion coefficients, densities) of both liquid water and ice polymorphs, as well as surface tensions, with accuracy very similar to TIP4P/2005, while performing very well for dielectric constants over a wide range of pressures and temperatures. On the other hand, TIP4P/2005 provides a better description of phase boundaries, including liquid-vapor and freezing transitions. Overall, these two models appear to perform almost identically on the conventional benchmark Vega and Abascal (2011) while significantly outperforming the OPC force field.
@@ VI Conclusions
Unfortunately, the good description of solution thermodynamics of TIP4P/ breaks down when one seeks an accurate account of transport properties, a problem that is also shared by the TIP4P/2005 model Blazquez, Conde, and Vega (2023). This points to the limitations of point charge models and the need to address the problem of polarization in a physically meaningful way for next generation force fields with enhanced transferability. Current advances in computer architecture might well remedy the computational overhead and make such models competitive in the very near future. In the meantime, TIP4P/2005 and TIP4P/ force fields appear as the two equally valid point charge models for use in routine calculations.
An intriguing question that remains unsolved is which of these two models is a better caricature of the actual charge distribution and local electric field created by a water molecule in the condensed phase. However, because this is actually a collective property that depends on many body interactions, it might simply be an ill-posed question. Be as it may, calculations of the dielectric constant of these models indicate that they predict a rather different Kirkwood factor, and accordingly, significantly different orientation correlations in the liquid phase. Measuring such correlations experimentally appears difficult, but it would allow us to discriminate between these two seemingly different empirical force fields. In the time being, our results illustrate that a model targeted to reproduce dielectric constants can perform as well as TIP4P/2005 for pure water, and allows for an improved transferability in solution. In this sense, it seems that a good parametrization can be afforded without invoking the otherwise appealing concept of charge scaling.
Acknowledgments
We would like to thank Eva G. Noya and Enrique Lomba for helpful discussions. LGM wishes to thank his thesis advisor Carlos Vega for his guidance and friendship during his PhD. ŁB would like to thank the Polish National Agency for Academic Exchange, under Grant No. BPN/BEK/2023/1/00006 Bekker 2023. CDAG acknowledges the predoctoral fellowship "Contratos Predoctorales de Personal Investigador en Formación" from the Universidad Complutense de Madrid (CT15/23) and additional funding from Ministerio de Ciencia, Innovación y Universidades (MICIU), under grant FPU22/00869. LGM would like to thank financial support from the Agencia Estatal de Investigación (Ministerio de Ciencia y Economía) under grant PID2023-151751NB-I00, funded by MCIN/AEI/10.13039/501100011033. We also benefited from generous allocation of computer time at the Academic Supercomputer Center (CI TASK) in Gdansk.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.